y"-4y'+9y=t,y(0)=0,y'(0)=1, find the transfer function


1
Expert's answer
2021-10-27T15:44:21-0400

Laplace transform of "y''-4y'+9y=t", is given as


"L( y''-4y'+9y)=L(t)"

"L(y'')=s^2L(y)-sy(0)-y'(0)"

"L(y')=sL(y)-y(0)"

"L(t)=\dfrac{1}{s^2}"


Putting the above values in the differential equations, we get,


"s^2L(y)-1-4sL(y)+9L(y)=\dfrac{1}{s^2}""L(s) (s^2-4s+9)=\frac{1}{s^2}+1"


or, 

"L(y)=\frac{1+s^2}{s^2(s^2-4s+9)}"